You roll two fair dice, a green one and a red one. (Round your answers to three decimal places.)
What is the probability of getting a sum of 11?

Answer by stanbon(75887)

You roll two fair dice, a green one and a red one. (Round your answers to three decimal places.)
What is the probability of getting a sum of 11?
-------
# of "11" combinations: 3
# of random combinations: 36
----
P(sum of 11) = 3/36 = 1/12
=====================
Cheers,
Stan H.
===========